# bibliography.bib

### v3

@@ 3 Estimating Markov chain convergence rates using Common Random Number Simulation
Fix $N\geq 1$. By the strong law of large numbers and since $E[d(X_{N},Y_{N})^{p}]<\infty$,
$\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{p}\overset{a.s.}{=}E[d(X_{N},Y_{N})^{p}]$. By Theorem 3.1, $W_{d,p}(\mathcal{L}(X_{N}),\pi)^{p}\leq\\ KE[d(X_{N},Y_{N})^{p}]$ and so equation 4 follows. ∎
The approach of Biswas and Mackey [biswas]
Algorithm 1 closely resembles Algorithm 1 of [biswas]. Biswas and Mackey show in [biswas] how CRN can be used to estimate the CUB estimator, which is defined as follows for a given $p\in[1,\infty)$,
$CUB_{M,N,T}=\left(\frac{1}{M(T-N)}\sum_{i=1}^{M}\sum_{n=N+1}^{T}d(X_{n,i},Y_{n,i})^{p}\right)^{1/p}$ (5)
By Corollary 3.2 of [biswas], under certain regularity conditions, the CUB estimator bounds the Wasserstein distance between the sample means,
$W_{d,p}\left(\mathcal{L}\left(\frac{1}{(T-N)}\sum_{n=N+1}^{T}X_{n,i}\right),\mathcal{L}\left(\frac{1}{(T-N)}\sum_{n=N+1}^{T}Y_{n,i}\right)\right)\leq\lim_{M\to\infty}CUB_{M,N,T}$
@@
$W_{d,p}(\mathcal{L}(X_{\infty}),\mathcal{L}(Y_{\infty}))\leq\lim_{M,N\to\infty,T-N\to\infty}CUB_{M,N,T}$
@@
In [biswas], the CUB estimator eventually bounds the Wasserstein distance between the corresponding stationary distributions, but does not tell us precisely when it will happen or by how much. That is, if the starting points of the two Markov chains are close in distance but are far from stationarity, the results from [biswas] would conclude, for some fixed iteration $N$, that the distance between the two chains is close (true), but it would be incorrect to state that the distance to stationarity is close. In comparison, Theorem 3.3 measures distance to stationarity for any initial distribution on $X_{0}$ and iteration $N\in\mathbb{N}$.
The approach of Johnson [crnconv]
@@
The motivation of using a rejection sampler comes from [crnconv], which provides a bound in total variation distance between a Markov chain and its corresponding stationary distribution using CRN simulation. The following is a simplification of the results with altered notation to match this paper. For complete details, refer to [crnconv]. For a discussion on when CRN is applicable within the setting of [crnconv], see [mcconvcomp].
@@
Let $\tau$ be a ‘stopping’ time, $\tau=\min\{n:d(X_{n},Y_{n})\leq\epsilon\}$. Theorem 2 of [crnconv] states that the total variation between two Markov chains $\{X_{n}\}_{n\geq 1}$, $\{Y_{n}\}_{n\geq 1}$ both with initial distribution $\nu$ and stationary distribution $\pi$ is as follows,
$||\pi-\mathcal{L}(X_{n})||_{TV}\leq\frac{2P(\tau\geq n)}{1-r(\pi,\nu)}+O(\epsilon)$
@@
The notation $O(\epsilon)$ means that for some fixed $\epsilon_{0}>0$, there exists an $M\in\mathbb{N}$ such that for any $\epsilon\in(0,\epsilon_{0})$, $||\pi-\mathcal{L}(X_{n})||_{TV}\leq\frac{2P(\tau\geq n)}{1-r(\pi,\nu)}+M\epsilon$.
@@
Both Theorem 3.3 and Theorem 2 of [crnconv] require calculating upper bounds on $r(\pi,\nu)$. However, Theorem 3.3 estimates an exact upper bound on the Wasserstein distance while Theorem 2 of [crnconv] estimates an upper bound on the total variation distance with an unknown component that converges to 00 as $\epsilon\to 0$.
@@
Applying total variation distance to the bound generated by Theorem 3.3 and taking $\epsilon=0$, we get a similar result to [crnconv], $||\pi-\mathcal{L}(X_{n})||_{TV}\leq\frac{P(\tau\geq n)}{1-r(\pi,\nu)}$. See Remark Remark for more details. Exact simulation of $d(X_{n},Y_{n})=0$ is not realistically attainable, however. This is why [crnconv] introduces the error $\epsilon$. This points to CRN simulation being a better fit for bounding Wasserstein distance compared to total variation distance.
@@ 4 Gibbs sampler for a variance component model
Consider a Gibbs sampler on the variance component model (see [varcomp]), which we will apply Theorem 3.3 to simulate convergence bounds on the Wasserstein distance. Convergence rate bounds on the total variation metric for the variance component model have already been discussed in Theorems 1 of [varcomp, Jones2004SufficientBF]. In Real Data Example 4.2, we compare our bound to the one generated in [simconv], which uses DnM arguments.
@@
[…]
@@
$f_{\nu}(\vec{\theta},v,w,\mu)=$ $f_{IG}(V\mid a_{1},b_{1}-1)f_{IG}(W\mid a_{2},b_{2}-1)f_{N}(\mu\mid a_{3},b_{3})\prod_{i=1}^{I}f_{N}(\theta_{i}\mid\bar{Y}_{i},W/J_{i})$ (7)
@@
Let $X_{n}=(\vec{\theta}_{n},V_{n},W_{n},\mu_{n})$ and $X^{\prime}_{n}=(\vec{\theta}^{\prime}_{n},V^{\prime}_{n},W^{\prime}_{n},\mu^{\prime}_{n})$ be two copies of the variance component Gibbs sampler initialized with $X^{\prime}_{0}\sim\nu$ where $f_{\nu}$ is defined in equation 7 and $X_{0}\perp\!\!\!\perp X^{\prime}_{0}$. Fix $p\in[1,\infty)$ and let $\pi$ be the corresponding stationary distribution. Then for prior parameters $b_{1},b_{2}>1$,
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq(KE[d(X_{n},X^{\prime}_{n})^{p}])^{1/p}$ (8)
@@
  • $K=\frac{C}{L}\left(\frac{1}{(V^{*})^{I/2}}e^{-\frac{1}{V^{*}}}\right)\left(\frac{1}{(W^{*})^{(\sum_{i=1}^{I}J_{i}-I)/2}}e^{-\frac{1}{W^{*}}}\right)$
  • $C=\left(\frac{b_{1}}{b_{1}-1}\right)^{a_{1}}\left(\frac{b_{2}}{b_{2}-1}\right)^{a_{2}}\frac{1}{(2\pi)^{(\sum_{i=1}^{I}J_{i})/2}}\frac{1}{(\prod_{i=1}^{I}J_{i})^{1/2}}$
  • $L=\exp\left[\int_{B}\log(g(\vec{\theta},v,w,\mu))d(\vec{\theta},V,W,\mu)\right]$ where $B\subset\mathbb{R}^{I}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\times\mathbb{R}$
  • $V^{*}=\frac{2}{I}$ and $W^{*}=\frac{2}{\sum_{i=1}^{I}J_{i}-I}$
@@
Note that in lemma 4.1 the value of $K$ still has an integral, which must be bounded from below, $L=\exp\left[\int_{B}\log(g(\vec{\theta},v,w,\mu))d(\vec{\theta},V,W,\mu)\right]$. To address this, we use the adaptIntegrate function in R and extract the lower bound.
@@
For this variance component model, upper bounds in total variation distance have been established. Theorem 1 of [varcomp] provides an upper bound on the total variation distance between the variance component Gibbs sampler model and its corresponding stationary distribution. A DnM condition to bound the distance in total variation is also presented in [Jones2004SufficientBF]. Discussions of convergence diagnostics on total variation distance for variations of this example can be found in [alicia, varcompgeom, roman, tan, varcompgeyer].
@@
Upper bounds in Wasserstein distance of the variance component model have also already been discussed. [wasmeth] analyzes the variance component model when the prior on $\mu$ is $\mu\propto 1$ instead of $\mu\sim N(a_{3},b_{3})$. In particular Proposition 25 of [wasmeth] states that if $\lim_{I\to\infty}J^{2}/(I^{3+\delta})=\infty$ for some $\delta>0$ and the model is properly specified (condition (E2) in [wasmeth]), then the geometric rate of convergence for this model goes to 00. Proposition 24 of [wasmeth] also bounds the total variation distance in terms of the Wasserstein distance. Discussions of convergence rates for the variance component model in Wasserstein distance are also presented in [davis, yang].

### v4

@@ 3 Estimating Markov chain convergence rates using CRN simulation
Since $\mathcal{L}(X_{N}),\mathcal{L}(Y_{N})\in P_{rs}(\mathcal{X})$, $E[d(X_{N},Y_{N})^{rs}]<\infty$ and so by the strong law of large numbers, $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{rs}\overset{a.s.}{=}E[d(X_{N},Y_{N})^{rs}]$. By Theorem 3.1, $W_{d,p}(\mathcal{L}(X_{N}),\pi)\leq K^{1/r}E[d(X_{N},Y_{N})^{rs}]^{\frac{1}{rs}}\overset{a.s.}{=}K^{1/r}\left(\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{rs}\right)^{\frac{1}{rs}}$
To optimize this bound we can choose $\nu$ that is close to $\pi$. If $K$ is very large, we can choose an $r\geq p$ such that $K^{1/r}$ is a reasonable value.
Next we state the confidence interval on our bound. We use this confidence interval in figures 1, 4, 5 and 6 and Table 6.
Lemma 3.3.
The $1-\alpha$ confidence interval on $K^{1/r}\left(\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{rs}\right)^{\frac{1}{rs}}$ is
$\left[K^{\frac{1}{r}}\left(m-z_{\alpha/2}\sigma_{m}/\sqrt{M}\right)^{\frac{1}{rs}},K^{\frac{1}{r}}\left(m+z_{\alpha/2}\sigma_{m}/\sqrt{M}\right)^{\frac{1}{rs}}\right]$
Where $m=\frac{1}{M}\sum_{i=1}^{M}d(X_{N,i},Y_{N,i})^{rs}$, $\sigma^{2}_{m}$ is the sample variance of $m$ and $\mathcal{L}(X_{N}),\mathcal{L}(Y_{N})\in P_{2rs}(\mathcal{X})$.
Proof.
Since $\mathcal{L}(X_{N}),\mathcal{L}(Y_{N})\in P_{2rs}(\mathcal{X})$, the variances are finite and we can apply the Central Limit Theorem. That is, $P\left(m-\frac{z_{\alpha/2}\sigma_{m}}{\sqrt{M}}\leq E[d(X_{n},Y_{n})^{rs}]\leq m+\frac{z_{\alpha/2}\sigma_{m}}{\sqrt{M}}\right)\approx 1-\alpha$. And so, $P\left(K^{\frac{1}{r}}\left(m-\frac{z_{\alpha/2}\sigma_{m}}{\sqrt{M}}\right)^{\frac{1}{rs}}\leq K^{\frac{1}{r}}E[d(X_{n},Y_{n})^{rs}]^{\frac{1}{rs}}\leq K^{\frac{1}{r}}\left(m+\frac{z_{\alpha/2}\sigma_{m}}{\sqrt{M}}\right)^{\frac{1}{rs}}\right)\approx 1-\alpha$. ∎
The approach of Biswas and Mackey [5]
Algorithm 1 closely resembles Algorithm 1 of [5]. Biswas and Mackey show in [5] how CRN simulation can be used to estimate the CUB estimator, which is defined as follows for a given $p\in[1,\infty)$,
$\text{CUB}_{M,N,T}=\left(\frac{1}{M(T-N)}\sum_{i=1}^{M}\sum_{n=N+1}^{T}d(X_{n,i},Y_{n,i})^{p}\right)^{1/p}$ (5)
By Corollary 3.2 of [5], under certain regularity conditions, the CUB estimator bounds the Wasserstein distance between the sample means,
$W_{d,p}\left(\mathcal{L}\left(\frac{1}{(T-N)}\sum_{n=N+1}^{T}X_{n,i}\right),\mathcal{L}\left(\frac{1}{(T-N)}\sum_{n=N+1}^{T}Y_{n,i}\right)\right)\leq\lim_{M\to\infty}\text{CUB}_{M,N,T}$
@@
$W_{d,p}(\mathcal{L}(X_{\infty}),\mathcal{L}(Y_{\infty}))\leq\lim_{M,N\to\infty,T-N\to\infty}\text{CUB}_{M,N,T}$
@@
In [5], the CUB estimator eventually bounds the Wasserstein distance between the corresponding stationary distributions, but does not tell us precisely when it will happen or by how much. That is, if the starting points of the two Markov chains are close in distance but are far from stationarity, the results from [5] would conclude, for some fixed iteration $N$, that the distance between the two chains is close (true), but it would be incorrect to state that the distance to stationarity is close. In comparison, Theorem 3.2 measures distance to stationarity for any initial distribution on $X_{0}$ and iteration $N\in\mathbb{N}$.
The approach of Johnson [31]
@@
The motivation of Theorem 3.2 comes from [31], which provides a bound in total variation distance between a Markov chain and its corresponding stationary distribution using CRN simulation. The following is a simplification of the results with altered notation to match this paper. For complete details, refer to [31]. For a discussion on when CRN is applicable within the setting of [31], see [12].
@@
Let $\tau$ be a ‘stopping’ time, $\tau=\min\{n:d(X_{n},Y_{n})\leq\epsilon\}$. Theorem 2 of [31] states that the total variation between two Markov chains $\{X_{n}\}_{n\geq 1}$, $\{Y_{n}\}_{n\geq 1}$ both with initial distribution $\nu$ and stationary distribution $\pi$ is as follows,
$\pi-\mathcal{L}(X_{n})||_{TV}\leq 2KP(\tau\geq n)+O(\epsilon)$
@@
Where $K=ess\sup_{z}\frac{f_{\pi}(z)}{f_{\nu}(z)}$. The notation $O(\epsilon)$ means that for some fixed $\epsilon_{0}>0$, there exists an $M\in\mathbb{N}$ such that for any $\epsilon\in(0,\epsilon_{0})$, $\pi-\mathcal{L}(X_{n})||_{TV}\leq\frac{2P(\tau\geq n)}{1-r(\pi,\nu)}+M\epsilon$.
@@
In [31], $K=\frac{1}{1-r(\pi,\nu)}$ where $r(\pi,\nu)$ is the rejection rate in a rejection sampler with stationary distribution $\pi$ and proposal distribution $\nu$ (See Definition 8.2). By Equation 8.2, $\frac{1}{1-r(\pi,\nu)}=ess\sup_{z}\frac{f_{\pi}(z)}{f_{\nu}(z)}$. Thus, both Theorem 3.2 (with $s_{1}=1,s_{2}=\infty$) and Theorem 2 of [31] require calculating upper bounds on $r(\pi,\nu)$. However, Theorem 3.2 estimates an exact upper bound on the Wasserstein distance while Theorem 2 of [31] estimates an upper bound on the total variation distance with an unknown component that converges to 0 as $\epsilon\to 0$.
@@
Applying total variation distance to the bound generated by Theorem 3.2 and taking $\epsilon=0$, we get a similar result to [31], $\pi-\mathcal{L}(X_{n})||_{TV}\leq KP(\tau\geq n)$. See Remark 2 for more details. In our simulations, exact simulation of $d(X_{n},Y_{n})=0$ was not realistically attainable, however. This is why [31] introduces the error $\epsilon$. This points to CRN simulation being a better fit for bounding Wasserstein distance compared to total variation distance.
4 Gibbs sampler for a model related to James-Stein estimators
Consider a Gibbs sampler for a model related to the James-Stein estimator, which we will apply Theorem 3.2 to simulate convergence bounds on the Wasserstein distance. The James-Stein estimator is similar to a variance component model (Example 5.1), but each $\theta_{i}$ has only one observation, $Y_{i}$. Its significance in the frequentist statistics literature comes from Stein’s paradox, which proves that for $\theta_{i}$ the estimator, $\left(1-\frac{I-2}{\sum_{j=1}^{I}Y_{j}^{2}}\right)_{+}Y_{i}$, has smaller risk than the intuitive estimator, $Y_{i}$, when $I\geq 3$ [58]. Here, we look at the same model setup, but through a Bayesian statistics lense. See [46] for more details on this example.
Example 4.1 (A model related to James-Stein estimators).
$Y_{i}|\theta_{i}\sim N(\theta_{i},V)$
The Gibbs sampler for a model related to James-Stein estimators.
Algorithm 2 shows how to apply a Gibbs sampler on this example.
Initialize $(\vec{\theta}_{0},\mu_{0},A_{0})$
for $n=1,\ldots,N$ do
  For each $i\in 1,\ldots,q$, generate $\theta_{n,i}|\mu_{n-1},A_{n-1},Y\sim N(\frac{Y_{i}A_{n-1}+\mu_{n-1}V}{A_{n-1}+V},\frac{VA_{n-1}}{V+A_{n-1}})$
end for
We will bound the Wasserstein distance between a Markov chain and its corresponding stationary distribution using Lemma 4.1.
Lemma 4.1.
Let $\{X_{n}\}_{n\geq 1}=(\vec{\theta}_{n},\mu_{n},A_{n})_{n\geq 1}$ and $\{X^{\prime}_{n}\}_{n\geq 1}=(\vec{\theta}^{\prime}_{n},\mu^{\prime}_{n},A^{\prime}_{n})_{n\geq 1}$ be two copies of the Markov chain initialized with $\vec{\theta}^{\prime}_{0}\sim N_{q}(\vec{Y},VI_{q})$, $A^{\prime}_{0}\sim IG(\alpha+(q-1)/2,\beta)$, $\mu^{\prime}_{0}\sim N(\bar{\theta}^{\prime}_{0},A^{\prime}_{0})$, and $X_{0}\perp\!\!\!\perp X^{\prime}_{0}$. Fix $p\in(1,\infty]$ and let $\pi$ be the corresponding stationary distribution. Then for $r\geq p$
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq\left(\frac{\Gamma(\alpha+(q-1)/2)(2\pi)^{1/2}}{\beta^{\alpha+(q-1)/2}}\frac{1}{L}E[d(X_{n},X^{\prime}_{n})^{r}]\right)^{1/r}$
Where $L=\int_{B}\frac{A^{-\alpha-1}e^{-\beta/A}}{(A+V)^{q/2}}e^{\frac{1}{2}\sum_{i=1}^{q}\left[\frac{(\frac{Y_{i}}{V}+\frac{\mu}{A})^{2}}{\frac{1}{V}+\frac{1}{A}}-\frac{Y_{i}^{2}}{V}-\frac{\mu^{2}}{A}\right]}\mathrm{d}(\mu,A)$ and $B\subset\mathbb{R}\times\mathbb{R}_{+}$ is a bounded set.
The proof is in section 8.4.
Real Data Example 4.1.
We apply our results to the batting averages ($Y$) for 18 ($q=18$) major league baseball players in 1970 [8, 27]. We want to find the estimated upper bound on the Wasserstein and total variation distance for a Gibbs sampler applied to the James-Stein estimator using the baseball data. We set the priors to $\alpha=0.01,\beta=2$ and $V=Var(Y)=0.00485$. The initial value of our Markov chains are independently drawn; $X_{0}=(\vec{\theta}_{0},\mu_{0},A_{0})=(100)^{q+2}$ and $X^{\prime}_{0}$ is distributed according to Lemma 4.1. Using Lemma 4.1, $K=5.9535$ and
[…]
holds almost surely.
[…]
Figure 1: The upper bound on the Wasserstein distance for the Gibbs sampler model related to James-Stein estimators. The estimate uses the mean of 1000 simulations and initial value $(\vec{\theta}_{0},\mu_{0},A_{0})=(100)^{q+2}$. The vertical axis is log scaled. $r=1$.
At iteration 9, the bound on the Wasserstein distance first hits below 0.01 and $W_{||\cdot||_{1},1}(\mathcal{L}(X_{9}),\pi)\leq 0.00073$.
Before we bound the Markov chain and the corresponding stationary distribution in total variation we introduce Theorem 4.8 of [51], which provides an upper bound in total variation based on the expected differences between the Markov chains through the constant $K_{TV}$. This theorem will also be used to generate the one-shot coupling bound.
Proposition 4.2 (Theorem 4.8 of [51]).
$\|\mathcal{L}(X_{n})-\mathcal{L}(X^{\prime}_{n})\|_{TV}$ $\leq K_{TV}E[|A_{0}-A^{\prime}_{0}|]\left(\frac{1}{q}\right)^{n-1}$ (7)
where $K_{TV}=\frac{(S/2)^{\frac{q-1}{2}}}{\Gamma(\frac{q-1}{2})}\left(\frac{S}{q+1}\right)^{-\frac{q-3}{2}}e^{-\frac{q+1}{2}}$ and $S=\sum_{i=1}^{q}(Y_{i}-\bar{Y})^{2}$.
Using Proposition 4.2, we get $K_{TV}=0.0047$. Therefore by Lemma 4.11 of [51],
[…]
So, by iteration 6, $\|\mathcal{L}(X_{6})-\pi\|_{TV}\leq 0.00866$.
We compare our bound to Proposition 4.2 and Theorem 4 of [46], which use one-shot coupling and DnM arguments, respectively, to generate a theoretical bound in total variation. The methodologies and assumptions are outlined in Table 1.
Method Common random number Drift and minorization One shot coupling
Source Lemma 4.1 Theorem 4 of [46] Proposition 4.2 and Proposition 6 of [40]
Prior $A\sim IG(0.01,2)$ and $f_{\mu}\propto 1$ $A\sim IG(-1,2)$ and $f_{\mu}\propto 1$ $f_{A}\propto 1$ and $f_{\mu}\propto 1$
Bound $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq$ $0.0282E[|X_{n}-X^{\prime}_{n}|]$ where $X_{n},X^{\prime}_{n}$ are simulated using CRN. $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq$ $0.967^{n}+(0.935)^{n}\times$ $(1.17+\sum_{i=1}^{q}(\theta_{i}-\bar{Y})^{2})$ $\|\mathcal{L}(X_{n})-\pi\|_{TV}\leq$ $\frac{0.0282}{17}E[|A_{0}-A_{1}|]\left(\frac{1}{18}\right)^{n}$
Note the minor variations in the examples, Proposition 4.2 assumes that the prior on $A$ is flat. In Theorem 4 of [46], the prior on $A\sim IG(-1,2)$ with the intent of generating a flat distribution. In our real data example we set the prior to $A\sim IG(0.01,2)$, which could approximate a flat distribution. We will assume the minor variations are comparable.
Table 3 and Figure 2 compare the bounds on total variation distance using three different methods: CRN, DnM, and one-shot coupling. The one-shot coupling bound indicates total variation distance of less than 0.01 by iteration 5. The DnM bound indicates total variation distance of less than 0.01 by iteration 249. Thus, the one-shot coupling bounds performs the best closely followed by the CRN bound and significantly better than the DnM bound.
Iteration CRN DnM One-Shot Coupling
1 56.2430 167,409.95 58.6595
2 336.0368 156,528.33 3.2589
3 20.0791 146,354.02 0.1811
4 1.4003 136,841.04 0.0101
5 0.0898 127,946.40 0.0005
6 0.0087 119,629.91 3.1049e-05
7 0.0007 111,853.99 1.7247e-06
8 6.0301e-05 104,583.51 9.5814e-08
9 3.4421e-06 97,785.60 5.3230e-09
10 2.6439e-07 91,429.56 2.9572e-10
249 1.6957e-307 0.0099 2.7286e-310
Comparison of bounds on $\mathcal{L}(\vec{\theta}_{n},\mu_{n},A_{n})-\pi||_{TV}$ Refer to caption
The bound generated from CRN simulation and one-shot coupling are expected to be very similar since to generate the geometric convergence rate of the one-shot coupling bound (the $1/18$ in table 1) we use CRN (see [28, 51, 52, 53]). This is confirmed by Figure 2 where the CRN bound and the one-shot bound are approximately parallel.
Finally, we compare our convergence rates with informal methods in Section 8.1.1. The Gelman-Rubin statistic is evaluated for each parameter. By iteration 5, the Gelman-Rubin statistic for all of the parameters graphed is less than 1.05. The traceplot paints a similar picture suggesting that all parameters have converged by iteration 6. Thus, the CRN bound indicating total variation distance convergence by iteration $6$, is tight.
@@ 5 Gibbs sampler for a variance component model
Consider a Gibbs sampler on the variance component model (see [48]), which we will apply Theorem 3.2 to simulate convergence bounds on the Wasserstein distance. Convergence rate bounds on the total variation metric for the variance component model have already been discussed in Theorems 1 of [48, 32, 11].
@@
$\mu$$\theta_{1}$$Y_{1,1},\ldots Y_{1,J_{1}}$$\theta_{2}$$Y_{2,1},\ldots Y_{2,J_{2}}\ldots$$\theta_{I}$$Y_{I,1},\ldots Y_{I,J_{I}}$
@@
$f_{\nu}(\vec{\theta},V,W,\mu)=$ $f_{IG}(V\mid 2a_{1}+1,2b_{1}-1)f_{IG}(W\mid 2a_{2}+1,2b_{2})f_{N}(\mu\mid a_{3},b_{3})\prod_{i=1}^{I}f_{N}\left(\theta_{i}\mid\bar{Y}_{i},\frac{W}{2J_{i}}\right)$ (9)
@@
Let $X_{n}=(\vec{\theta}_{n},V_{n},W_{n},\mu_{n})$ and $X^{\prime}_{n}=(\vec{\theta}^{\prime}_{n},V^{\prime}_{n},W^{\prime}_{n},\mu^{\prime}_{n})$ be two copies of the variance component Gibbs sampler initialized with $X^{\prime}_{0}\sim\nu$ where $f_{\nu}$ is defined in Equation 9 and $X_{0}\perp\!\!\!\perp X^{\prime}_{0}$. Fix $p\in[1,\infty)$ and let $\pi$ be the corresponding stationary distribution. Then for $r\geq p,b_{1}>1/2$,
$W_{d,p}(\mathcal{L}(X_{n}),\pi)\leq K^{1/r}E[d(X_{n},X^{\prime}_{n})^{2r}]^{\frac{1}{2r}}$ (10)
@@
  • $K=\frac{1}{L}C_{1}C_{2}^{1/2}$
  • $L=\int_{B}(b_{1}^{*})^{-a_{1}^{*}}e^{-\frac{(\mu-a_{3})^{2}}{2b_{3}}}\left(b_{2}^{*}\right)^{-(a_{2}^{*})}\mathrm{d}(\vec{\theta},\mu)$ where $B\subset\mathbb{R}^{I+1}$
  • $C_{1}=\frac{\Gamma(a_{1})}{\Gamma(a_{1}^{*})b_{1}^{a_{1}}}\frac{\Gamma(a_{2})}{\Gamma(a_{2}^{*})b_{2}^{a_{2}}}(2\pi)^{\sum_{i=1}^{I}J_{i}/2+(I+1)/2}b_{3}^{1/2}$
  • $C_{2}=\frac{b_{1}^{2a_{1}}}{\Gamma(a_{1})^{2}}\frac{\Gamma(2a_{1}+1)}{(2b_{1}-1)^{2a_{1}+1}}\frac{b_{2}^{2a_{2}}}{\Gamma(a_{2})^{2}}\frac{\Gamma(2a_{2}+1)}{(2b_{2})^{2a_{2}+1}}\frac{\Gamma(I/2-1)}{(\prod_{i=1}^{I}J_{i})^{1/2}2^{I+\sum_{i=1}^{I}J_{i}}\pi^{T+I/2}}\frac{\Gamma(T-1)}{S^{T-1}}$
  • $a_{1}^{*}=a_{1}+I/2$ and $b_{1}^{*}=b_{1}+\frac{\sum_{i=1}^{I}(\theta_{i}-\mu)^{2}}{2}$
  • $a_{2}^{*}=a_{2}+\sum_{i=1}^{I}J_{i}/2$ and $b_{2}^{*}=b_{2}+\sum_{i=1}^{I}\sum_{j=1}^{J_{i}}\frac{(Y_{ij}-\theta_{i})^{2}}{2}$
  • $S=\sum_{i=1}^{I}(\sum_{j=1}^{J_{i}}Y_{ij}^{2}-J_{i}(\bar{Y}_{i})^{2})$ and $T=\sum_{i=1}^{I}J_{i}-I/2$
@@
Note that in Lemma 5.1 the value of $K$ still has an integral, which must be bounded from below, $L=\int_{B}g(\vec{\theta},v,w,\mu)\mathrm{d}(\vec{\theta},V,W,\mu)$. To address this, we use the adaptIntegrate function in R.
For this variance component model, upper bounds in total variation distance have been established. In Real Data Example 5.1, we compare our bounds in Wasserstein distance to the total variation bounds generated in [32].
@@
[48] also provides an upper bound on the total variation distance between the variance component Gibbs sampler model and its corresponding stationary distribution (See Theorem 1 of [48]). However the constants required to obtain an explicit upper bound are difficult to calculate (see Remark 6 of [48]). An upper bound in total variation distance for this example is also established in [11]. Discussions of convergence diagnostics on total variation distance for variations of this example can be found in [30, 1, 45, 54, 23].
@@
Upper bounds in Wasserstein distance of the variance component model have also already been discussed. [40] analyzes the variance component model when the prior on $\mu$ is $\mu\propto 1$ instead of $\mu\sim N(a_{3},b_{3})$. In particular Proposition 25 of [40] states that if $\lim_{I\to\infty}J^{2}/(I^{3+\delta})=\infty$ for some $\delta>0$ and the model is properly specified (condition (E2) in [40]), then the geometric rate of convergence for this model goes to 0. Proposition 24 of [40] also bounds the total variation distance in terms of the Wasserstein distance. Discussions of convergence rates for the variance component model in Wasserstein distance are also presented in [14, 59].
Numerical Example 5.1.
We are going to compare our convergence bounds to the bounds from [32]. The data for this example is simulated in [32] and summarised in Table 4. [32] generates an upper bound on the total variation distance between a Markov chain and its corresponding stationary distribution using a DnM bound stated in [47] (see Theorem 12 of [47] and Theorem 3.1 of [32]). The hyperparameters for the DnM bound are defined in Table 5.
Cell 1 2 3 4 5
$\bar{Y}_{i}$ -0.80247 -1.0014 -0.69090 -1.1413 -1.0125
Table 4: Summary of simulated data used in [32] (see their Table 1) and in Numerical Example 5.1. $I=5$, $J=10$, $\sum_{i=1}^{5}\sum_{j=1}^{10}(y_{ij}-\bar{y}_{i})^{2}=32.990$
Hyperparameter $a_{1}$ $b_{1}$ $a_{2}$ $b_{2}$ $a_{3}$ $b_{3}$
From [32] and this paper 2.5 1 1 1 $\bar{Y}$ 1
Table 5: Hyperparameters used in Numerical Example 5.1 and Table 2 of [32]. $\bar{Y}=\frac{1}{IJ}\sum_{i=1}^{I}\sum_{j=1}^{J}Y_{ij}$
Applying Lemma 5.1, $K=48.7818$ and the following holds for $r=1$,
[…]
The initial value of $X_{0}=(\vec{\theta}_{0},V_{0},W_{0},\mu_{0})$ is $V_{0}\sim IG(a_{1},b_{1})$, $W_{0}\sim IG(a_{2},b_{2})$ and $(\vec{\theta}_{0},\mu_{0})=$ the initial value defined in [32], Remark 4.1 of [32]. The initial value of $X^{\prime}_{0}$ satisfies Equation 9.
[…]
Figure 4: A Wasserstein distance upper bound on the Gibbs sampler for a variance component model (Numerical Example 5.1) using CRN. The estimate uses the mean of 1000 simulations with initial values of $\mu_{0},\vec{\theta}_{0}$ satisfying the initial values defined in [32] (see their Remark 4.1). The vertical axis is a log scale.
[…]
1 397506.3 0 3075.601 55941.57
2 6.0002 112.7212 119.4924 125.8998
3 0.1835 17.3169 20.8980 23.9495
4 0.0076 3.2230 4.2734 5.1123
5 0.0004 0.7855 0.9132 1.0251
6 2.4539e-05 0.1837 0.2417 0.2881
7 1.4015e-06 0.0441 0.0578 0.0687
8 1.0485e-07 0.0114 0.0158 0.0192
9 5.4138e-09 0.0027 0.0036 0.0043
Table 6: Wasserstein bound estimates and confidence intervals for the variance component model, Numerical Example 5.1.
The CRN bound indicates Wasserstein distance of less than 0.0099 by iteration 9 ($W_{||\cdot||_{1},1}(\mathcal{L}(X_{9}),\pi)\\ \leq 0.00359$). The bound for [32] indicates total variation distance of less than 0.00999 by iteration 34153415 ($\mathcal{L}(X_{3415})-\pi||_{TV}\leq 0.00999$) (See Table 3 of [32]). Note that the CRN bound is measured in Wasserstein distance while the bound in [32] is measured in total variation distance. The large disparity between the DnM and the CRN bound is consistent with the conclusions of [39] that say that the convergence rate bounds based on single-step DnM tend to be overly conservative.
Finally we compare our convergence rates with informal methods in Section 8.1.2. The traceplot (Figure 8) suggests that convergence of $\mu$, $V$, and $W$ is nearly immediate. The Gelman-Rubin statistic is evaluated at different iterations in Table 9. By iteration 10, the Gelman-Rubin statistic is close to or less than 1.05 for $V,W$, and $\mu$.